A right triangle is a triangle with one angle measuring 90°. In a right triangle, the sides are related by Pythagorean Theorem, , where c is the hypotenuse (the side opposite the 90° angle). Find the hypotenuse when the other 2 sides’ measurements are 6 feet and 8 feet.
Answer by Nate(3500)

The Pythagorean Theorem is c^2=a^2+b^2 for a right triangle.
you can use tangent, cosin, or sin to find the hypotenuse, but we will use the easy way. "a" is the measurement of one of the sides, while "b" is the measurement of the other side.
c^2=a^2+b^2
c^2=(8)^2+(6)^2
c^2=64+36
c^2=100
c=10
The hypotenuse is 10 feet!